# Learning Multiple Initial Solutions to Optimization Problems

### v1

@@ Abstract
Sequentially solving similar optimization problems under strict runtime constraints is essential for many applications, such as robot control, autonomous driving, and portfolio management. The performance of local optimization methods in these settings is sensitive to the initial solution: poor initialization can lead to slow convergence or suboptimal solutions. To address this challenge, we propose learning to predict multiple diverse initial solutions given parameters that define the problem instance. We introduce two strategies for utilizing multiple initial solutions: (i) a single-optimizer approach, where the most promising initial solution is chosen using a selection function, and (ii) a multiple-optimizers approach, where several optimizers, potentially run in parallel, are each initialized with a different solution, with the best solution chosen afterward. We validate our method on three optimal control benchmark tasks: cart-pole, reacher, and autonomous driving, using different optimizers: DDP, MPPI, and iLQR. We find significant and consistent improvement with our method across all evaluation settings and demonstrate that it efficiently scales with the number of initial solutions required. The code is available at https://github.com/EladSharony/miso.
@@ 1 Introduction
Figure 1: As opposed to previous works that predict a single initial solution, MISO trains a single neural network to predict multiple initial solutions. We use them to either initialize a single optimizer or jointly initialize multiple optimizers.
@@ 3 Initializing optimizers
$\bm{x}^{\star}(\bm{\psi})=\arg\min_{\bm{x}}J(\bm{x};\bm{\psi})\quad\mathrm{s.t.}\quad\bm{g}(\bm{x};\bm{\psi})\leq\bm{0};\quad\bm{h}(\bm{x};\bm{\psi})=\bm{0},$ (1)
@@ 4.1 Multi-output predictor
Our multi-output predictor is a neural network that takes the problem instance, $\bm{\psi}$, as input and outputs $K$ initial solutions for the optimization problem,
@@ 4.2 Optimization with multiple initial solutions
In the single-optimizer setting we run a single instance of the optimizer with the most promising initial solution, $\bm{\hat{x}^{\star}}=\bm{\mathrm{Opt}}(J,\bm{\psi},t_{\textrm{limit}};\bm{\hat{x}}^{\mathrm{init}})$. We introduce a selection function, $\bm{\Lambda}$, which, given a set of candidate solutions and the problem instance $\bm{\psi}$, returns the most promising candidate, $\bm{\hat{x}}^{\mathrm{init}}=\bm{\Lambda}(\{\bm{\hat{x}}_{k}^{\mathrm{init}}\}_{k=1}^{K},\bm{\psi})$. A reasonable choice for $\bm{\Lambda}$ used in our experiments is selecting the candidate that minimizes the objective function the optimizer aims to minimize, i.e., $\bm{\Lambda}:=\arg\min_{k}J(\bm{\hat{x}}_{k}^{\mathrm{init}};\bm{\psi})$. Other possibilities include risk measures, metrics based on performance stability, robustness, exploration, or domain-specific metrics that align with the objectives of the overall task.
@@
Our framework can be trivially generalized to allow a different number of optimizers and initial solution predictions, as well as using a heterogeneous set of optimization methods. Further, to maintain performance guarantees, one may include the default, e.g., warm-start, solution as one of the considered initial solutions, which ensures that even with poor predictions, the final solution quality does not degrade.
@@ 4.3 Training strategies
$\mathcal{L}_{\mathrm{MISO-PD}}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})+\alpha_{K}\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{\mathrm{PD},k}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star}),$
@@
$\mathcal{L}_{\mathrm{MISO-WTA}}=\min_{k}\{\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})\}.$
@@
$\mathcal{L}_{\mathrm{MISO-MIX}}=\min_{k}\left\{\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})+\alpha_{K}\Phi\left(\mathcal{L}_{\mathrm{PD},k}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})\right)\right\},$
@@ 4.4 Illustrative example
Figure 2: Top: The one-dimensional cost function $c(x)$ with global minima at $\bm{\mathrm{A}}$ and $\bm{\mathrm{C}}$ and a local minimum at $\bm{\mathrm{B}}$. Bottom: Predicted initial solution for different methods, demonstrating why explicitly promoting multimodality is important.
@@ 5 Experimental setup
We employ two evaluation modes. (i) One-off, where the optimization task is treated as an isolated problem with the objective of finding the minimum of a given function. This mode serves as the default configuration for training neural networks, where data is replayed to the model, and the optimizer’s solution is recorded but not executed. Methods are assessed by the mean cost of the optimizer’s output over problem instances. (ii) Sequential, which involves solving a series of related optimization problems, executing each proposed solution, and starting the subsequent optimization from the resulting state. This setting simulates real-world conditions where the optimizer continuously interacts with a dynamic environment in a closed loop. We evaluate performance by taking the mean cost over problems in a sequence, and then the mean over sequences.
@@
To generate the training data, we first create a set of problem instances by sequentially executing the optimizer initialized with the default warm-start strategy. The problem instances are then fed again to an “oracle” version of the optimizer with a significantly increased runtime limit, and the resulting solutions are recorded. After training, evaluation is done on a separate unseen set of problem instances. All experiments are conducted on an Intel Core i9-13900KF CPU and an NVIDIA RTX 4090 GPU. Further implementation details, including hyperparameters and training procedures, are in Appendix A.4 and Appendix A.3.
@@ 6 Results
Figure 4: Mean Cost of the autonomous driving environment with varying values of $K$. The shaded regions around each curve indicate the standard error of the mean.
@@
Single optimizer
. In the single-optimizer setting, Table 1, we first observe that even one learned initialization outperforms heuristic solutions (regression vs. warm-start), in almost all settings, and in particular in the most challenging autonomous driving task. We then examine the impact of generating multiple initial solutions. Perturbations-based methods show some improvement over their single-initialization counterparts in most cases, and ensembles of independently learned models perform consistently better than single models. Finally, our proposed multi-output methods demonstrate substantial improvements over all baselines because they can learn to predict diverse multimodal initial solutions. Specifically, MISO winner-takes-all or MISO mix achieve the lowest mean costs across all tasks. Considering the pairwise distance term alone proves insufficient to ensure adequate diversity, whereas incorporating it with MISO winner-takes-all often boosts performance, yet, its effectiveness varies, which underscores the challenge of selecting optimal hyperparameters. As expected, improvements are consistently larger in the more important sequential optimization setting, where errors over time compound.
Multiple optimizers
. When considering the multiple-optimizers setting, we observe the same trend. Learning-based methods outperform heuristic ones, and multi-output approaches yield further enhancements. As expected, the use of multiple optimizers leads to consistently better results compared to the single-optimizer setting due to increased exploration of the solution space.
@@
Figure 5 (left) depicts the optimizer’s output trajectories with different initial solutions for the autonomous driving task. In this scenario, the high-level planner abruptly alters the reference path, which could happen, e.g., because of a newly detected pedestrian. The change in reference path makes the previous solution (warm-start) a poor initialization, and the optimizer converges to a local minimum that minimizes control effort but is far from the desired path. Regression and model ensemble also fail to predict a good initial solution. In contrast, MISO winner-takes-all adapts to this sudden reference change and closely follows the reference path. Figure 5 (right) depicts MISO’s initial solutions for the cart-pole task. The different outputs capture different modes of the solution space (moving upright, moving with left swing, moving with right swing), showing MISO’s ability to generate diverse and multimodal solutions.
@@
Figure 5: Left: Each method adaptation when the high-level planner abruptly modifies the reference path. Right: Multiple predicted trajectories of MISO winner-takes-all.
@@ 7 Conclusions and Future Work
Other extensions of our approach include probabilistic modeling, e.g., Gaussian mixture models, variational autoencoders, or diffusion models; however, preventing mode collapse and promoting diversity would remain a challenge. Future work may explore alternative selection functions, such as risk measures or criteria based on stability, robustness, exploration, or other domain-specific metrics; as well as using a heterogeneous set of parallel optimizers. Finally, we are excited about various possible applications in optimal control and beyond, where sequences of similar optimization problems need to be solved, for example, localization and mapping in robotics, financial optimization, traffic routing optimization, or even training neural networks with different initial weights, e.g., for meta-learning, or scene representation learning with Neural Radiance Fields or 3D Gaussian splatting.
Ethics statement.
Our work is concerned with a general class of optimization problems that do not raise particular ethical considerations.
@@ A.3 Network Architecture and Training Details
In all experiments, the training data is generated by (1) unrolling an optimizer using a warm-start initialization policy and recording its inputs and outputs, and (2) replaying the same scenarios using an oracle optimizer—essentially the same optimizer with enhanced capabilities, such as more optimization steps or additional sampled trajectories—and logging its inputs and outputs. The resulting mapping may be seen as a filter, refining (near-)optimal initial solutions into optimal ones.
@@ A.7 State Loss
An additional challenge in learning control policies is addressing compounding errors—small inaccuracies in the predicted control trajectory $\bm{\tau}_{u}$ that, when unrolled, cause significant deviations in the state trajectory $\bm{\tau}_{x}$. Even if most elements $\bm{\tau}_{u}$ are accurate, errors in the initial steps can cause the state $\bm{\tau}_{x}$ to drift, leading to further divergence as the system evolves.
To mitigate compounding errors, we introduce a regression loss not only over the control trajectory $\bm{\tau}_{u}$ but also over the resulting state trajectory $\bm{\tau}_{x}$. In a supervised learning setting, this requires a model of the system dynamics, which can either be known or learned. As we show in Table 6, incorporating state trajectory loss helps mitigate these types of error accumulation and improve long-horizon trajectory accuracy. More specifically, we see that (1) as the prediction horizon increases, from $H=9$ (Cart-pole) to $H=40$ (Driving), so does the difference between using and not using state loss, (2) The gap between One-Off and Sequential also increases thus we do not generalize as well, (3) For the single-output regression model, the difference is even greater.

### v2

@@ Abstract
Sequentially solving similar optimization problems under strict runtime constraints is essential for many applications, such as robot control, autonomous driving, and portfolio management. The performance of local optimization methods in these settings is sensitive to the initial solution: poor initialization can lead to slow convergence or suboptimal solutions. To address this challenge, we propose learning to predict multiple diverse initial solutions given parameters that define the problem instance. We introduce two strategies for utilizing multiple initial solutions: (i) a single-optimizer approach, where the most promising initial solution is chosen using a selection function, and (ii) a multiple-optimizers approach, where several optimizers, potentially run in parallel, are each initialized with a different solution, with the best solution chosen afterward. Notably, by including a default initialization among predicted ones, the cost of the final output is guaranteed to be equal or lower than with the default initialization. We validate our method on three optimal control benchmark tasks: cart-pole, reacher, and autonomous driving, using different optimizers: DDP, MPPI, and iLQR. We find significant and consistent improvement with our method across all evaluation settings and demonstrate that it efficiently scales with the number of initial solutions required. The code is available at MISO.
Figure 1: As opposed to previous works (top) that predict a single initial solution, MISO trains a single neural network to predict multiple initial solutions. We use them to either initialize a single optimizer (middle) or jointly initialize multiple optimizers (bottom).
@@ 1 Introduction
Notably, any existing initialization strategy can be combined with MISO, by simply including the existing initial solution among the predicted ones; and by design, MISO is guaranteed to be equal or better than the default initialization.
@@ 3 Initializing optimizers
$\bm{x}^{\star}(\bm{\psi})=\arg\min_{\bm{x}}J(\bm{x};\bm{\psi})\quad\text{s.t.}\quad\begin{aligned} \bm{g}(\bm{x};\bm{\psi})&\leq\bm{0},\\ \bm{h}(\bm{x};\bm{\psi})&=\bm{0}.\end{aligned}$ (1)
@@ 4.1 Multi-output predictor
Our multi-output predictor is a standard Transformer model–chosen for its simplicity and clarity (see Appendix A.13 for a discussion on alternative architectures). It takes the problem instance, $\bm{\psi}$, as input and outputs $K$ initial solutions for the optimization problem,
@@ 4.2 Optimization with multiple initial solutions
In the single-optimizer setting we run a single instance of the optimizer with the most promising initial solution, $\bm{\hat{x}^{\star}}=\bm{\mathrm{Opt}}(J,\bm{\psi},t_{\textrm{limit}};\bm{\hat{x}}^{\mathrm{init}})$. We introduce a selection function, $\bm{\Lambda}$, which, given a set of candidate solutions and the problem instance $\bm{\psi}$, returns the most promising candidate, $\bm{\hat{x}}^{\mathrm{init}}=\bm{\Lambda}(\{\bm{\hat{x}}_{k}^{\mathrm{init}}\}_{k=1}^{K},\bm{\psi})$. A reasonable choice for $\bm{\Lambda}$ used in our experiments is selecting the candidate that minimizes the objective function the optimizer aims to minimize, i.e., $\bm{\Lambda}:=\arg\min_{k}J(\bm{\hat{x}}_{k}^{\mathrm{init}};\bm{\psi})$. However, other criteria–such as robustness, constraint satisfaction, or domain-specific requirements–may be more appropriate in certain scenarios; see Appendix A.10 for details and examples.
@@
Guarantees.
Our framework can be trivially generalized to allow a different number of optimizers and initial solution predictions, as well as using a heterogeneous set of optimization methods. To maintain performance guarantees, one may include traditional initialization methods, such as warm-start heuristics, as part of the set of initializations. MISO is guaranteed to improve over the existing default strategy by design. In the single-optimizer setting the best initial solution is always equal or better than the default according to the selection function $\bm{\Lambda}$; and in the multiple-optimizer setting the final solution is equal or better than using only the default initialization.
@@ 4.3 Training strategies
$\mathcal{L}_{\mathrm{PD}}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})+\alpha_{K}\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{\mathrm{PD},k}(\bm{\hat{x}}_{k}^{\mathrm{init}}),$
@@
$\mathcal{L}_{\mathrm{WTA}}=\min_{k}\{\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})\}.$
@@
$\mathcal{L}_{\mathrm{MIX}}=\min_{k}\left\{\mathcal{L}_{\mathrm{reg}}(\bm{\hat{x}}_{k}^{\mathrm{init}},\bm{x}^{\star})+\alpha_{K}\Phi\left(\mathcal{L}_{\mathrm{PD},k}(\bm{\hat{x}}_{k}^{\mathrm{init}})\right)\right\},$
@@ 4.4 Illustrative example
Figure 2: Left: The one-dimensional cost function $c(x)$ with global minima at $\bm{\mathrm{A}}$ and $\bm{\mathrm{C}}$ and a local minimum at $\bm{\mathrm{B}}$. Right: Predicted initial solution for different methods, demonstrating why explicitly promoting multimodality is important.
@@ 5 Experimental setup
We employ two evaluation modes. (i) One-off, where the optimization task is treated as an isolated problem with the objective of finding the minimum of a given function. This mode serves as the default configuration for training neural networks, where data is replayed to the model, and the optimizer’s solution is recorded but not executed. Methods are assessed by the mean cost of the optimizer’s output over problem instances. (ii) Sequential, which involves solving a series of related optimization problems, executing each proposed solution, and starting the subsequent optimization from the resulting state. This setting simulates real-world conditions where the optimizer continuously interacts with a dynamic environment in a closed loop. Astute readers may notice parallels to the open-loop/closed-loop paradigms in control theory; these connections are discussed in greater depth in Appendix A.11 and Appendix A.12
We evaluate performance by taking the mean cost over problems in a sequence, and then the mean over sequences.
@@
To generate the training data, we first create a set of problem instances by sequentially executing the optimizer initialized with the default warm-start strategy. The problem instances are then fed again to an ”oracle” version of the optimizer with a significantly increased runtime limit, and the resulting solutions are recorded. After training, evaluation is done on a separate unseen set of problem instances. All experiments are conducted on an Intel Core i9-13900KF CPU and an NVIDIA RTX 4090 GPU. Further implementation details, including hyperparameters and training procedures, are in Appendix A.3 and Appendix A.4.
@@ 6 Results
Figure 4: Mean cost of the driving task (Sequential Optimization) for varying $K$ values. The shaded regions indicate the standard error.
Figure 5: Single Optimizer for Driving (left) and Cart-Pole (right). On the left, we show each method’s adaptation when the high-level planner abruptly modifies the reference path. On the right, we illustrate multiple trajectories predicted by MISO winner-takes-all.
@@
Single optimizer.
In the single-optimizer setting, Table 1, we first observe that even one learned initialization outperforms heuristic solutions (regression vs. warm-start), in almost all settings, and in particular in the most challenging autonomous driving task. We then examine the impact of generating multiple initial solutions. Perturbations-based methods show some improvement over their single-initialization counterparts in most cases, and ensembles of independently learned models perform consistently better than single models. Finally, our proposed multi-output methods demonstrate substantial improvements over all baselines because they can learn to predict diverse multimodal initial solutions. Specifically, MISO winner-takes-all or MISO mix achieve the lowest mean costs across all tasks. Considering the pairwise distance term alone proves insufficient to ensure adequate diversity, whereas incorporating it with MISO winner-takes-all often boosts performance, yet, its effectiveness varies, which underscores the challenge of selecting optimal hyperparameters. As expected, improvements are consistently larger in the more important sequential optimization setting, where errors over time compound.
Multiple optimizers.
When considering the multiple-optimizers setting, we observe the same trend. Learning-based methods outperform heuristic ones, and multi-output approaches yield further enhancements. As expected, the use of multiple optimizers leads to consistently better results compared to the single-optimizer setting due to increased exploration of the solution space.
@@
Performance guarantees.
To evaluate the guarantees discussed in Sect. 4.2, we added the warm-start initial solution to the set of candidates of MISO winner-takes-all and MISO mix. We observed in all problem instances the cost of the best initial solution to be equal or lower than the cost of the warm start in the single optimizer setting; and the cost of the final solution to be lower or equal than for the warm start in the multiple optimizer setting. The mean costs remained mostly similar to MISO, and improved slightly in some cases. Detailed results are in Appendix A.15.
@@
Figure 5 (left) depicts the optimizer’s output trajectories with different initial solutions for the autonomous driving task. In this scenario, the high-level planner abruptly alters the reference path, which could happen, e.g., because of a newly detected pedestrian. The change in reference path makes the previous solution (warm-start) a poor initialization, and the optimizer converges to a local minimum that minimizes control effort but is far from the desired path. Regression and model ensemble also fail to predict a good initial solution. In contrast, MISO winner-takes-all adapts to this sudden reference change and closely follows the reference path. Figure 5 (right) depicts MISO’s initial solutions for the cart-pole task. The different outputs capture different modes of the solution space (maintaining balance while moving, swinging leftward, and swinging rightward), showing MISO’s ability to generate diverse and multimodal solutions.
@@ 7 Conclusions and Future Work
Other extensions of our approach include probabilistic modeling, e.g., Gaussian mixture models, variational autoencoders, or diffusion models; however, preventing mode collapse and promoting diversity, and inference overhead (see Appendix A.14), would remain a challenge. Future work may explore alternative selection functions, such as risk measures or criteria based on stability, robustness, exploration, or other domain-specific metrics; as well as using a heterogeneous set of parallel optimizers. Finally, we are excited about various possible applications in optimal control and beyond, where sequences of similar optimization problems need to be solved, for example, localization and mapping in robotics, financial optimization, traffic routing optimization, or even training neural networks with different initial weights, e.g., for meta-learning.
Impact statement.
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3 Network Architecture and Training Details
In all experiments, the training data is generated by (1) unrolling an optimizer using a warm-start initialization policy and recording its inputs and outputs, and (2) replaying the same scenarios using an oracle optimizer—essentially the same optimizer with enhanced capabilities, such as more optimization steps or additional sampled trajectories—and logging its inputs and outputs. The resulting mapping may be seen as a filter, refining (near-)optimal initial solutions into optimal ones. For each task, 500,000 instances were collected.
@@ A.7 State Loss
An additional challenge in learning control policies is addressing compounding errors—small inaccuracies in the predicted control trajectory $\bm{\tau}_{u}$ that, when unrolled, cause significant deviations in the state trajectory $\bm{\tau}_{x}$. Even if most elements of $\bm{\tau}_{u}$ are accurate, errors in the initial steps can cause the state $\bm{\tau}_{x}$ to drift, leading to further divergence as the system evolves.
To mitigate compounding errors, we introduce a regression loss not only over the control trajectory $\bm{\tau}_{u}$ but also over the resulting state trajectory $\bm{\tau}_{x}$. In a supervised learning setting, this requires a model of the system dynamics, which can either be known or learned.
Let $\hat{\bm{\tau}}_{u}=\{\hat{\bm{u}}_{t}\}_{t=0}^{H-1}$ denote the predicted control trajectory, and $\bm{\tau}_{u}^{\star}=\{\bm{u}^{\star}_{t}\}_{t=0}^{H-1}$ denote the target control trajectory. Similarly, let $\hat{\bm{\tau}}_{x}=\{\hat{\bm{x}}_{t}\}_{t=1}^{H}$ be the predicted state trajectory obtained by unrolling the predicted controls through the system dynamics starting from the initial state $\bm{x}_{0}$, i.e.,
$\hat{\bm{x}}_{t+1}=f(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t}),\quad\text{with }\hat{\bm{x}}_{0}=\bm{x}_{0},$ (2)
and let $\bm{\tau}_{x}^{\star}=\{\bm{x}^{\star}_{t}\}_{t=1}^{H}$ be the target state trajectory.
We define the control loss as
$\mathcal{L}_{\text{control}}=\frac{1}{H}\sum_{t=0}^{H-1}\left\|\hat{\bm{u}}_{t}-\bm{u}^{\star}_{t}\right\|^{2},$ (3)
and the state loss as
$\mathcal{L}_{\text{state}}=\frac{1}{H}\sum_{t=1}^{H}\left\|\hat{\bm{x}}_{t}-\bm{x}^{\star}_{t}\right\|^{2}.$ (4)
Our total loss function combines these two components:
$\mathcal{L}=\mathcal{L}_{\text{control}}+\lambda\,\mathcal{L}_{\text{state}},$ (5)
where $\lambda$ is a weighting factor that balances the contributions of the control loss and the state loss.
By incorporating the state loss $\mathcal{L}_{\text{state}}$, we encourage the predicted control trajectory to produce a state trajectory that remains close to the target state trajectory, thereby mitigating compounding errors during rollout.
As we show in Table 6, incorporating state trajectory loss helps mitigate these types of error accumulation and improve long-horizon trajectory accuracy. More specifically, we see that (1) as the prediction horizon increases, from $H=9$ (Cart-pole) to $H=40$ (Driving), so does the difference between using and not using state loss, (2) The gap between One-Off and Sequential also increases thus we do not generalize as well, (3) For the single-output regression model, the difference is even greater.
@@ A.9 Mean Cost Sequential Optimization
A.10 Selection Function $\Lambda$
In our framework, the selection function $\Lambda$ plays a critical role in choosing the most promising initial solution from the set of candidates predicted by our model. While using the objective function $J$ as $\Lambda$ is a natural and effective choice, alternative choices for $\Lambda$ can be advantageous in certain scenarios.
A.10.1 Alternative Selection Criteria
Constraint Satisfaction: In some applications, especially those that are safety-critical, it is essential to ensure that certain constraints are satisfied by the initial solution, even if it means accepting a higher value of $J$. In such cases, $\Lambda$ can be designed to prioritize solutions that satisfy these constraints. For example:
$\Lambda(\hat{\mathbf{x}}_{k}^{\text{init}},\psi)=J(\hat{\mathbf{x}}_{k}^{\text{init}},\psi)+\beta\,C(\hat{\mathbf{x}}_{k}^{\text{init}},\psi),$ (6)
where $C(\hat{\mathbf{x}}_{k}^{\text{init}},\psi)$ measures the degree of constraint violation, and $\beta$ is a weighting factor that penalizes constraint violations.
Robustness Measures: $\Lambda$ can incorporate robustness criteria, selecting initial solutions less sensitive to model uncertainties or external disturbances. For instance, it could favor solutions that maintain performance across various scenarios.
Contextual Adaptation: The selection function can adapt based on the problem instance $\psi$. For example, in varying environmental conditions, $\Lambda$ could prioritize more conservative or aggressive solutions depending on the context or operational requirements.
A.10.2 Learning the Selection Function
Instead of hand-crafting $\Lambda$, it can be learned from data. One approach is to model $\Lambda$ as a parameterized function, such as a neural network, and train it jointly with the predictor model or separately. The learning objective could be to maximize the overall performance of the optimizer when initialized with the selected solutions, potentially incorporating criteria like safety, robustness, or energy efficiency.
A.10.3 Examples
Safety-Critical Control: In autonomous driving, safety constraints such as maintaining a safe distance from obstacles are crucial. $\Lambda$ can prioritize trajectories that ensure safety over those that simply minimize time or fuel consumption. For example, it can assign infinite cost to any solution violating safety constraints, effectively excluding unsafe options.
Adaptive Behavior: In robotics, $\Lambda$ can select initial solutions that favor energy efficiency when the robot’s battery is low or prioritize speed when tasks are time-sensitive. By incorporating the robot’s current state or mission objectives into $\Lambda$, the system can adapt its behavior accordingly.
A.11 Evaluation Modes: One-off and Sequential
In our experiments, we assess the performance of the methods using two evaluation modes:
One-off Evaluation: In the one-off evaluation, problem instances are uniformly sampled from the evaluation dataset (disjoint from the training dataset). Each method is tested on the same set of independently sampled instances, ensuring a fair comparison across methods. The optimizer solves each problem instance independently, without any interaction with the environment or influence from previous solutions. This evaluation mode focuses on the optimizer’s ability to find high-quality solutions for individual problems in isolation.
Sequential Evaluation: In sequential evaluation, the optimizer interacts with the environment across a series of time steps. Starting from an initial state sampled from the evaluation dataset (disjoint from the training dataset), at each time step the optimizer adjusts its decisions based on the evolving state. This mode evaluates the optimizer’s performance in a dynamic, real-time setting, highlighting its ability to manage evolving states and adapt over time.
A.12 Sequential vs. Closed-loop and One-off vs. Open-loop
In control settings, the terms sequential and closed-loop evaluations are often used interchangeably. However, in the context of general optimization problems, the notion of ”closed-loop” may not always be applicable, as there may be no dynamic environment or feedback mechanism involved. Therefore, we adopt more general terminology—referring to sequential evaluations—to encompass scenarios where decisions are made in a sequence but without necessarily involving feedback from an environment.
On the other hand, the distinction between one-off and open-loop evaluations is subtle yet significant. In a one-off evaluation, we assess the optimizer’s performance on individual problem instances without any interaction with an environment. This means the optimizer solves a static problem, and we can directly compare different methods on the same set of instances. In contrast, open-loop control involves sequentially executing actions in an environment without feedback.
A.13 Alternative Neural Backbone Architectures
We chose a standard Transformer backbone primarily for its simplicity and clarity. Our focus is not on comparing different sequence-based architectures, e.g., Action Chunking Transformers Zhao et al. (2023), LSTM variants, or hierarchical models, but on introducing a multi-output framework that can be integrated into a wide range of neural backbones to generate multiple diverse initial solutions. Consequently, replacing the Transformer with, say, an Action Chunking Transformer or another advanced architecture would still benefit from the same multi-output concept and diversity-promoting objectives, with minimal implementation changes.
A.14 Sampling-Based Architectures
Methods that rely on sampling multiple solutions at inference time—such as diffusion models—are particularly ill-suited for real-time settings with strict runtime constraints. Generating $K$ distinct samples typically entails either multiple iterative steps per sample or repeated forward passes, causing inference time to scale poorly with $K$. This overhead is incompatible with scenarios like autonomous driving, where any significant increase in computation time can undermine real-time performance. For instance, a Diffusion Policy Chi et al. (2024) took [^1] Mean over 1000 runs using Cadene et al. (2024) implementation, on an NVIDIA RTX 4090 GPU. $378.50$ milliseconds to produce a single sequence for the autonomous driving task (length $40$ and dimension $2$), compared to $K$ >>1 sequences in just $0.85$ milliseconds using a multi-output regressor.
A.15 Experiments on combining MISO with warm-start
In Table 7 and Table 8 we report mean cost results for combining MISO predicted initial solution candidates with the heuristic warm-start initial solution. Specifically MISO denotes a setting where the warm-start is added as an extra candidate. Results show that MISO consistently improves over warm-start. Further, including warm-start as one of the candidate initial solutions can sometimes lead to additional improvements (MISO vs. MISO).
Table 7: Results for combining MISO with warm-start for the single optimizer setting.
One-Off Optimization Sequential Optimization
Method $K$ Reacher Cart-pole Driving Reacher Cart-pole Driving
Single Optimizer
MISO winner-takes-all 32 13.36  $\pm$0.88 10.47  $\pm$0.77 29.95  $\pm$2.04 2.70  $\pm$0.21 0.82  $\pm$0.05 29.62  $\pm$1.98
MISO mix 32 12.74  $\pm$0.86 10.47  $\pm$0.77 33.19  $\pm$2.39 2.30  $\pm$0.19 0.77  $\pm$0.04 33.82  $\pm$2.48
Table 8: Results for combining MISO with warm-start for the multiple optimizers setting.
One-Off Optimization Sequential Optimization
Method $K$ Reacher Cart-pole Driving Reacher Cart-pole Driving
Multiple Optimizers
MISO winner-takes-all 32 13.34  $\pm$0.88 10.29  $\pm$0.76 30.62  $\pm$2.24 2.19  $\pm$0.17 0.76  $\pm$0.05 29.71  $\pm$1.96
MISO mix 32 12.72  $\pm$0.86 10.28  $\pm$0.76 33.26  $\pm$2.29 1.53  $\pm$0.14 0.63  $\pm$0.02 33.61  $\pm$2.36